OpenStudy (thehumantorch10):

Solve for a. a+1/4-a+2/6+a+3/9=2 a.0 b.7 c.9 d.63 Medal and Fan for answer and explanation

OpenStudy (lynfran):

\[a+\frac{ 1 }{ 4 }-a+\frac{ 2 }{ 6 }+a+\frac{ 3 }{ 9 }=2\]\[a-a+a=2-\frac{ 1 }{ 4 }-\frac{ 2 }{ 6 }-\frac{ 3 }{ 9}\]\[a=???\]

OpenStudy (thehumantorch10):

hmmm.. i dont understand the whole lcd thing.

OpenStudy (lynfran):

ok is that the exact problem that i wrote about? cause im not getting any of ur options ...the LCM is ...ok what are multiples of 4..they are 4, 8, 12... multiples of 6 are 6, 12,18.... multiples of 9 are, 9,18, 27 ,36... and you look to see what is the lowest common multiple in (4,6 and 9)...example...multiples of 2 and 4 multiple of 2 ...2,4,6,8 multiple of 4....4,8,12,16 etc the Lowest common Multiple between 2 and 4 is "4"...

OpenStudy (thehumantorch10):

ok......

OpenStudy (thehumantorch10):

hmmmm i dont know... the choices are 0,7,9 and 63. i just double-checked

OpenStudy (thehumantorch10):

im supposed to be solving rational equations.. the calculator must have solved it wrong.

OpenStudy (lynfran):

ok..

OpenStudy (anonymous):

ans is 9.

OpenStudy (anonymous):

want to how? just reply to me.